OpenStudy (anonymous):

Solve for x. 9^(x+8) = 17^(7x) using log base 10?

OpenStudy (anonymous):

a start is \[log(9^{x+8})=log(17^{7x})\] \[(x+8)log(9)=7xlog(17)\]

OpenStudy (anonymous):

remember that \[log(9)\] and \[log(17)\] are just numbers (constants) so it is algebra from there on in. multiply out etc to solve for x.

OpenStudy (anonymous):

\[xlog(9) + 8log(9)=7xlog(17)\] \[8log(9) = 7xlog(17)-xlog(9)\] \[8log(9)=x(7log(17)-log(9))\] \[x=\frac{8log(9)}{7log(17)-log(9)}\]

OpenStudy (anonymous):

and how do you find like what log(9) is ?

OpenStudy (anonymous):

use your calculator at this point